# Imperceptible Rhythm Backdoor Attacks: Exploring Rhythm Transformation for Embedding Undetectable Vulnerabilities on Speech Recognition

### v2

@@ 4.5 Ablation Study
Effects of Poisoning Number. We replace the poisoning rate with the poisoning number because our proposed trigger is effective and needs a very low number of poisoning samples. The poisoning rate corresponding to 50 samples is $0.22\%$. As shown in Table 7, the ASR increases with the increase in the number of poisoning number in general. It is indicated that in cases of low poisoning rates, compared to squeezing.
Effects of Rhythm Transformation Ratio. The best effectiveness is shown in the set of $0.5$ and $1.5$ rhythm transformation ratios. With squeezing, a relatively low poisoning number of 150 can achieve an excellent attack success rate of $99.97\%$. With stretching, a relatively low poisoning number of 150 can achieve an excellent attack success rate of $99.10\%$. The squeezing ASR values under each poisoned number are obviously higher than the stretching ones. This also indicates that squeezing behaves better than stretching.
Table 7: Ablation study. The ASR with different hyper-parameters
@@ 5 Conclusion
The paper proposed a speech backdoor attack method, combining mainly VAD, RSRT and neural vocoder. This method achieves very high ASR while maintaining a very low poisoning rate. The proposed trigger can avoid two kinds of main detection by speaker verification system and automatic speech recognition and gains excellent stealthiness and speech quality. The experiments demonstrate the excellent performance of efficiency and stealthiness of speech backdoor attacks with our method. We think that changing the rhythm or speech components of speech is an exploratory new approach to speech backdoor attacks.

### v3

@@ 4.5 Ablation Study
Effects of Poisoning Number: We replace the poisoning rate with the poisoning number because our proposed trigger is effective and needs a very low number of poisoning samples. The poisoning rate corresponding to 50 samples is $0.22\%$. As shown in Table 7, the ASR increases with the increase in the number of poisoning number in general. It is indicated that in cases of low poisoning rates, compared to squeezing.
Effects of Rhythm Transformation Ratio: The best effectiveness is shown in the set of $0.5$ and $1.5$ rhythm transformation ratios. With squeezing, a relatively low poisoning number of 150 can achieve an excellent attack success rate of $99.97\%$. With stretching, a relatively low poisoning number of 150 can achieve a superb attack success rate of $99.10\%$. The squeezing ASR values under each poisoned number are obviously higher than the stretching ones. This also indicates that squeezing behaves better than stretching.
Table 7: Ablation study: the ASR with different hyper-parameters
@@ 5 Conclusion
The paper proposed a speech backdoor attack method called RSRT, mainly combining VAD, RSRT, and neural vocoder. This method achieves very high ASR while maintaining a shallow poisoning rate. The proposed trigger can avoid two kinds of main detection by speaker verification system and automatic speech recognition and gains excellent stealthiness and speech quality. The experiments demonstrate the superb performance of efficiency and stealthiness of speech backdoor attacks with our method. We think that changing the rhythm or some speech components of speech is an exploratory new approach to speech backdoor attacks. In future research, the RSRT method can be applied to different languages and speaker scenarios to verify its generalization and cross-linguistic effectiveness. Considering that current speech recognition and speaker verification systems are continuously improving in terms of security and detection capabilities, future research could also focus on how to maintain a high level of stealthiness in more complex attack environments while developing more intelligent adversarial trigger generation strategies.